The difference of two numbers is 16 and the sum of their squares is 130. Find the two numbers.
I got -9 and -7, but one should be positive and I don't know what I have done wrong?

Answer by lwsshak3(11628)

The difference of two numbers is 16 and the sum of their squares is 130. Find the two numbers.
let x=one of the two numbers
let y=other number
x-y=16
x=y+16
..
x^2+y^2=130
(y+16)^2+y^2=130
y^2+32y+256+y^2=130
2y^2+32y+126=0
y^2+16y+63=0
(y+7)(y+9)=0
y=-7
x=y+16=9
or
y=-9
x=y+16=7
two sets of answers:
-7 & 9 or -9 & 7